Calculate the weight of Fe3O4 in 100.0g Fe2O3.

Given: molar mass of Fe2O3 =159.69 g/mol Molar mass of Fe3O4 = 231.54 g/mol Hint: you need two conversion factors
Answer #1

Given : grams of Fe2O3 = 100 g

molar mass of Fe2O3 = 159.69 g/mol

moles of Fe2O3 = mass in grams/ molar mass = 100/159.69 = 0.626 mole

the we will do our calculations with 0.626 mole of Fe2O3

The reaction is not given, we are assuming its thermal decomposition of Fe2O3(iron(III) oxide ) which takes place at a temperature around 1200 degree Celsius to make Fe3O4

6Fe2O3 ---> 4Fe3O4 + O2

From the above reaction:

molar ratio of Fe2O3 : Fe3O4 is 6 :4

this means

6 moles of Fe2O3 makes = 4 mol of Fe3O4

1 mol of Fe2O3 makes = 4/6 moles of Fe3O4

0.626 mol of Fe2O3 makes = 4/6 ] x 0.626 moles of Fe3O4 = 0.417 moles of Fe3O4

molar mass of Fe3O4 = 231.54 g/mol

mass in grams of Fe3O4 = moles x molar mass

= 0.417 x 231.54 = 96.66 g of Fe3O4